# Characterisations of Sobolev spaces and constant functions over metric spaces

### v2

@@ Abstract.
The “defect”, that the fractional Sobolev norm $\|f\|_{\dot{W}^{s,p}}$, for $s\in(0,1)$, fails to converge to the first-order Sobolev norm $\|f\|_{\dot{W}^{1,p}}$ as $s\to 1$, is a source of two types of substitute results: (1) characterisations of constant functions by integral conditions obtained by naïve substitution of $s=1$ into the formula of $\|f\|_{\dot{W}^{s,p}}$, and (2) derivative-free characterisations of $\dot{W}^{1,p}$ by suitable modifications of the said formula.
In this paper, we obtain extensions of both types of results for the versions of these function spaces over any complete doubling metric measure space $(X,\rho,\mu)$ supporting a $(1,p)$-Poincaré inequality. As a key tool of independent potential, we introduce a new “macroscopic” Poincaré inequality, whose right-hand side has oscillations of the same form as the left-hand side, but at a smaller macroscopic scale $r\in(0,R)$.
Besides intrinsic interest, these results are motivated by applications to quantitative compactness properties of commutators $[T,f]$ of singular integrals and pointwise multipliers. With pivotal use of the present results, a characterisation of commutator mapping properties, over the same class of general domains $(X,\rho,\mu)$, is obtained in a companion paper.
@@
The authors were supported by the Research Council of Finland through projects 346314 (T.H.), 360184 (R.K.), and 364208 (T.H.).
@@ 1. Introduction
The well-known “defect” [8], that the fractional Sobolev semi-norm
$\|f\|_{\dot{W}^{s,p}(\mathbb{R}^{d})}:=\Big[\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\Big(\frac{|f(x)-f(y)|}{|x-y|^{s}}\Big)^{p}\frac{\,\mathrm{d}x\,\mathrm{d}y}{|x-y|^{d}}\Big]^{\frac{1}{p}},\quad\begin{cases}s\in(0,1),\\ p\in(1,\infty),\end{cases}$ (1.1)
does not converge to the first-order Sobolev semi-norm $\|f\|_{\dot{W}^{1,p}(\mathbb{R}^{d})}:=\|\nabla f\|_{L^{p}(\mathbb{R}^{d})}$ as $s\to 1$, is a source of interesting results of two main types:
@@
    Showing that the finiteness of naïve versions of the limit, or the integral expression at $s=1$, actually force the function $f$ to be a constant, one obtains integral criteria to recognise constant functions (as in [10]).
@@
    Cleverly modifying the limit (as in [8]), or the integral expression at $s=1$ (as in [12]), one can “rectify the situation” [8] and recover the $\dot{W}^{1,p}$ norm, thus obtaining derivative-free characterisations of Sobolev spaces.
The theme of this paper is results of both types for Sobolev spaces over a metric measure space $(X,\rho,\mu)$ in place of $\mathbb{R}^{d}$. Despite previous literature on this theme at large (some of which we will shortly recall), our characterisations are not only new but also in demand for applications that we sketch in Section 1.B.2.
@@
for all $x\in X$ and $r>0$. For the main results, we also assume that $(X,\rho)$ is a complete metric space and supports the $(1,p)$-Poincaré inequality (Definition 2.2). We will informally refer to these as the “usual assumptions” (cf. [26, 27]).
In some examples, but not in our main results, the space is also assumed to have lower dimension $d$ or upper dimension $D$, by which we mean that
$\Big(\frac{R}{r}\Big)^{d}\lesssim\frac{V(x,R)}{V(x,r)}\lesssim\Big(\frac{R}{r}\Big)^{D}$ (1.2)
@@
for all $x\in X$ and $0<r\leq R<\infty$. In particular, a space is called Ahlfors $d$-regular, if (1.2) holds with $d=D$, or equivalently, $V(x,r)\approx r^{d}$. Every doubling space has upper dimension $D\leq\log_{2}c_{\mu}$, but possibly much lower than this simple upper bound. While not a property of every doubling space, a lower dimension $d>0$ readily follows from the “usual assumptions”: these imply that $X$ is connected by [3, Proposition 4.2], and a connected doubling space has $d>0$ by [3, Corollary 3.8].
1.A. Recognising constants
Starting with theme (1), a prototypical result [10, Proposition 2] states that the naïve substitution of $s=1$ into (1.1) is a characterisation of constants:
$\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\Big(\frac{|f(x)-f(y)|}{|x-y|}\Big)^{p}\frac{\,\mathrm{d}x\,\mathrm{d}y}{|x-y|^{d}}<\infty\quad\Leftrightarrow\quad f\equiv\text{const a.e.}$ (1.3)
While different extensions of (1.1) have been considered over $(X,\rho,\mu)$, it seems that the following fractional Sobolev (or Besov, hence the notation) norm is the most relevant one for many purposes [4, 5, 20, 21]:
$\|f\|_{\dot{B}^{s}_{p,p}(\mu)}:=\Big[\iint_{X\times X}\Big(\frac{|f(x)-f(y)|}{\rho(x,y)^{s}}\Big)^{p}\frac{\,\mathrm{d}\mu(x)\,\mathrm{d}\mu(y)}{V(x,y)}\Big]^{\frac{1}{p}},$ (1.4)
where
$V(x,y):=\mu(B(x,\rho(x,y)))\approx V(y,x),$
and the last equivalence is an easy consequence of the doubling condition.
Despite several extensions [6, 7, 25, 43, 44] of the classical result (1.3) to variants of the norm (1.4) and/or different assumptions on the underlying space $(X,\rho,\mu)$, the most natural generalisation (in our opinion)
$\|f\|_{\dot{B}^{1}_{p,p}(\mu)}<\infty\quad\Leftrightarrow\quad f\equiv\text{const a.e.},$ (1.5)
under the “usual assumptions” (in particular, without assuming Ahlfors regularity present in [6, 7, 43]) seems to have become available only very recently, in [35, Corollary 1.8].
Unaware of [35] when starting this project, we provide another proof of (1.5) under the same “usual assumptions”. Even after [35], we believe that our alternative proof has independent interest, being relatively direct in using only the Besov norms as defined in (1.4), while the proof of [35, Corollary 1.8]—quoting [1, Theorem 5.1], [27, Theorem 10.5.2], and [35, Theorem 1.6]—passes through the theory of Korevaar–Schoen spaces $KS^{1,p}$. In fact, we prove a slightly stronger local version of (1.5) in Proposition 4.1, as well as a variant in Corollary 4.3, which is relevant for an application that we discuss in some more detail in Section 1.B.2 below.
1.B. Characterisations of Sobolev spaces
Turning to our main theme (2), a sample of results along this extensive line of research in $\mathbb{R}^{d}$ is as follows:
  1. (a)
    As a prototypical case among several variants, [8] proved that
    $\|f\|_{\dot{W}^{1,p}}=c_{p,d}\lim_{s\to 1}(1-s)^{\frac{1}{p}}\|f\|_{\dot{W}^{s,p}},$
    which also shows the blow-up of the $\dot{W}^{s,p}$ norms in the lack of a renormalisation. Further formulas for $\|f\|_{\dot{W}^{1,p}}$ in terms of limits or supremums of fractional Sobolev type quantities are contained in [9, 40, 41]; see also [42] for a survey.
  2. (b)
    Replacing the $L^{p}$ norm in (1.1) by the weak $L^{p,\infty}$ norm when $s=1$, [12] found their surprising formula for the Sobolev norm
    $\|Q_{\frac{d}{p}}f\|_{L^{p,\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d})}\approx\|f\|_{\dot{W}^{1,p}(\mathbb{R}^{d})}=c_{p,d}\lim_{\lambda\to 0}\lambda|\{(x,y):Q_{\frac{d}{p}}(x,y)>\lambda\}|^{\frac{1}{p}}.$
    where $Q_{\gamma}f(x,y):=\frac{|f(x)-f(y)|}{|x-y|^{1+\gamma}}$. Originally a priori relations for test functions $f\in C_{c}^{\infty}(\mathbb{R}^{d})$ in [12], these were extended to all $f\in L^{1}_{\operatorname{loc}}(\mathbb{R}^{d})$ (and other values of $\gamma$) in [11], achieving a characterisation of the space $\dot{W}^{1,p}(\mathbb{R}^{d})$, rather than just the norm of test functions.
  3. (c)
    Another surprising formula using a weak $L^{p,\infty}$ norm is due to [19]. Denoting
    $m_{f}(x,t):=\fint_{B(x,t)}|f-\langle f\rangle_{B(x,t)}|,\qquad\,\mathrm{d}\nu_{p}(x,t):=\frac{\,\mathrm{d}x\,\mathrm{d}t}{t^{p+1}},$
    for $(x,t)\in\mathbb{R}^{d}\times(0,\infty)$, [19] showed that
    $\|m_{f}\|_{L^{p,\infty}(\nu_{p})}\approx\|f\|_{\dot{W}^{1,p}(\mathbb{R}^{d})}=c_{p,d}\lim_{\lambda\to 0}\lambda\cdot\nu_{p}(\{m_{f}>\lambda\})^{\frac{1}{p}}.$
    In the special case $p=d$, a variant of the norm comparison (but not the limit relation) was already contained in [14] and, with a different proof and under the a priori assumption that $f\in L^{\infty}(\mathbb{R}^{d})$, in [39]. All these results were motivated by applications to commutators $[T,f]$ discussed in Section 1.B.2.
1.B.1. Sobolev spaces over metric measure spaces
A potential use of such derivative-free characterisations of $\dot{W}^{1,p}$ is that they suggest meaningful definitions of a first-order Sobolev norm in any metric measure space $(X,\rho,\mu)$, consistent (by the results just quoted) with the usual definition on $\mathbb{R}^{d}$. However, since other notions of first-order Sobolev spaces have already proven to be useful in various contexts beyond $\mathbb{R}^{d}$, one would also like to have consistency in a broader class of domains $(X,\rho,\mu)$. We will explicitly work with the Hajłasz–Sobolev space $\dot{M}^{1,p}$ of [22] (see Definition 2.8), since this seems the most convenient choice for our considerations. However, under the “usual assumptions”, this is known to coincide with several other definitions (see Remark 2.11 for details), thus providing an essentially canonical first-order Sobolev space, against which other candidate definitions should be compared. In this direction, the following results have been obtained:
  1. (A)
    Under the “usual assumptions”, several of the limit formulas in (a) have been extended in [16, 36, 37].
  2. (B)
    Under the “usual assumptions” on $(X,\rho,\mu)$, the “surprising formulas” of [12] in (b) where extended by [15] to a special class of test functions on $X$. Under somewhat different assumptions on $(X,\rho,\mu)$, similar formulas were obtained by [24] for all Lipschitz functions of bounded support. Thus, in both cases, one obtains an equivalent expression for the Sobolev norm of test functions, as in [12], but not (yet) a characterisation of the full Sobolev space as in [11].
  3. (C)
    The norm comparison in (c) has been extended to the case of general stratified Lie groups (or Carnot groups) $\mathbb{G}$ in place of $\mathbb{R}^{d}$ in [38, Theorem A.3]. The key elements of the Euclidean argument of [19] are still available in this setting, making their adaptation in [38, pp. 110–112] a rather quick task. On the other hand, a variant of the earlier $p=d$ result of [14] has been obtained by [6, Theorem 1.4] under the “usual assumptions” plus the additional conditions that $(X,\rho,\mu)$ is a compact, Ahlfors regular space of homogeneous dimension $d$.
Our goal in this paper is an extension of (c) to general metric measure spaces under the “usual assumptions”, which requires significant departures from the Euclidean argument [19] and its Carnot group extension [38]. Compared to [6], besides dispensing with the compactness and Ahlfors regularity assumptions, our substitute of $\mathbb{R}^{d}\times(0,\infty)$ from [19] will be simply $X\times(0,\infty)$, instead of a hyperbolic filling of $X$, as in [6], which requires more effort to describe. This is our main result:
1.6 Theorem.
Let $(X,\rho,\mu)$ be a complete doubling metric measure space supporting a $(1,p)$-Poincaré inequality for some $p\in(1,\infty)$. Let $f\in L^{1}_{\operatorname{loc}}(\mu)$. Then $f$ belongs to the homogeneous Hajłasz–Sobolev space $\dot{M}^{1,p}(\mu)$ (Definition 2.8) if and only if $m_{f}\in L^{p,\infty}(\nu_{p})$, where $m_{f}$ and $\nu_{p}$ are the function and the measure on $X\times(0,\infty)$ defined by
In this case, we have
$\|f\|_{\dot{M}^{1,p}(\mu)}$ $\approx\|m_{f}\|_{L^{p,\infty}(\nu_{p})}:=\sup_{\kappa>0}\kappa\cdot\nu_{p}(\{m_{f}>\kappa\})^{1/p}$ (1.8)
$\approx\liminf_{\kappa\to 0}\kappa\cdot\nu_{p}(\{m_{f}>\kappa\})^{1/p}.$ (1.9)
The norm comparison (1.8) is a complete extension of the corresponding results in $\mathbb{R}^{d}$ [19] and Carnot groups [38]; see Examples 2.10 and 2.12 for the coincidence of $\dot{M}^{1,p}(\mu)$ with the usual Sobolev spaces in the respective settings. On the other hand, our limit relation (1.9) is somewhat weaker than the corresponding identity proved by [19] and recorded in (c). In the lack of the symmetries available in $\mathbb{R}^{d}$, we believe that (1.9) is the best substitute that one can expect in our generality.
1.B.2. Connection to commutators
Rather than being just another characterisation Sobolev spaces over metric measure spaces, the particular type of characterisation of Theorem 1.6 (just like its Euclidean and Carnot group predecessors) is motivated by its role in the characterisation of quantitative compactness properties of commutators $[T,f]$ of singular integrals $T$ and pointwise multipliers at a critical end-point case. On $\mathbb{R}^{d}$, for a class of operators $T$ containing, and modelled after, the classical Riesz transforms $\partial_{j}(-\Delta)^{-\frac{1}{2}}$, it is known that
$\|[T,f]\|_{S^{d,\infty}(L^{2}(\mathbb{R}^{d}))}\approx\|f\|_{\operatorname{Osc}^{d,\infty}(\mathbb{R}^{d})}\approx\|m_{f}\|_{L^{d,\infty}(\nu_{d})}\approx\|f\|_{\dot{W}^{1,d}(\mathbb{R}^{d})};$ (1.10)
here $S^{d,\infty}(L^{2}(\mathbb{R}^{d}))$ is the weak Schatten class of compact operators on $L^{2}(\mathbb{R}^{d})$, whose $d$th singular value decays at the rate $O(n^{-\frac{1}{d}})$, and $\operatorname{Osc}^{d,\infty}(\mathbb{R}^{d})$ is an oscillatory norm introduced by [45], who also proved the first $\approx$” in (1.10). The second $\approx$” is elementary, while the last one is (c) with $p=d$, the case already due to [14]. Another proof of $\|[T,f]\|_{S^{d,\infty}(L^{2}(\mathbb{R}^{d}))}\approx\|f\|_{\dot{W}^{1,d}(\mathbb{R}^{d})}$ without the intermediate steps of (1.10) is due to [39], under the additional assumption that $f\in L^{\infty}(\mathbb{R}^{d})$ and using the theory of pseudodifferential operators, which seems unlikely to extend to very general $(X,\rho,\mu)$.
With the goal of proving a version of this commutator result in reasonably general spaces, the approach of [45] and [14, 19] via (1.10) seemed like a more promising path. In the companion paper [29], the first two $\approx$” in (1.10) are extended to $(X,\rho,\mu)$ in place of $\mathbb{R}^{d}$; curiously, this part assumes Alhfors $d$-regularity in the second $\approx$” but no Poincaré at all. To complete the commutator theorem, it then remains to extend the last $\approx$” in (1.10), which is precisely (c) from above. For the full result, it is essential that we obtain the norm comparison for all $f\in L^{1}_{\operatorname{loc}}(\mu)$, instead of just test functions (as in the extensions of (b) described in (B)): we want that the Schatten property of the commutator implies the membership of $f$ in the Sobolev space, without imposing any a priori conditions of $f$. We achieve this in our Theorem 1.6, providing decisive input for the proof of the main theorem of [29].
The characterisation of constants also plays a role in [29] in showing that
$b,T]\in S^{p}(L^{2}(\mu))\quad\Rightarrow\quad b=\text{const}$
for certain values of $p$, extending a classical cut-off phenomenon [30] from $\mathbb{R}^{d}$ to metric measure spaces $(X,\rho,\mu)$ under the “usual assumptions”.
1.C. Plan of the paper and an overview of the argument
In the following Section 2, we collect some preliminaries, including the definitions omitted in this Introduction. In Section 3, we present our key new tool of the “macroscopic” Poincaré inequality in Theorem 3.4. With this result at hand, we obtain a rather quick proof of the characterisation of constant functions in Section 4.
The rest of the paper is dedicated to the proof of our Main Theorem 1.6. The rough outline of our proof is adapted from the Euclidean argument of [19]. In particular, the lower bound of the Sobolev norm (i.e., $\gtrsim$” in (1.8)) is obtained in Section 5 by a straightforward adaptation of [19]. However, the other direction is much more challenging, and we need to develop substantial new methods. Besides the macroscopic Poincaré inequality, which also features in this proof, we establish some delicate comparison of maximal and mean oscillations at the infinitesimal scale in Section 6, to substitute tools like Taylor expansions, which are not available in our setting. We only obtain this comparison for Lipschitz functions, and to lift this a priori assumption, we need to be able to approximate general functions by Lipschitz functions with respect to the “exotic” norm $f\mapsto\|m_{f}\|_{L^{p,\infty}(\nu_{p})}$ featuring in the desired characterisation. (This approximation follows easily from Theorem 1.6 a posteriori, once we know that the exotic norm is equivalent to $\|f\|_{\dot{M}^{1,p}(\mu)}$; but we need it a priori, in order to prove this equivalence!) This is accomplished in Section 7 with the help of the macroscopic Poincaré inequality, this time serving to substitute some convolution tricks, which are easy in $\mathbb{R}^{d}$, but unavailable in our setting. After these preparations, we can complete the proof of the upper bound of the Sobolev norm ($\lesssim$” in (1.8)) in Section 8 and the limit relation (1.9) in Section 9.
2. Preliminaries

### v3

@@ Abstract.
In a doubling metric measure space $(X,\rho,\mu)$ supporting a Poincaré inequality, we give a new characterisation of first-order Sobolev spaces by mean oscillations, and extend previous characterisations of constant functions in terms of the finiteness of certain integrals through a new approach. As a key tool of independent potential, we introduce a novel “macroscopic” Poincaré inequality, whose right-hand side has oscillations of the same form as the left-hand side, but at a smaller macroscopic scale $r\in(0,R)$.
Besides intrinsic interest, these results are motivated by applications to quantitative compactness properties of commutators $[f,T]$ of pointwise multipliers and singular integrals. With pivotal use of the present results, a characterisation of commutator mapping properties, over the same class of general domains $(X,\rho,\mu)$, is obtained in a companion paper.
@@
The authors were supported by the Research Council of Finland through projects 346314 (T.H.), 360184 (R.K.), 364208 (T.H.), and 371637 (T.H.).
@@ 1. Introduction
This investigation arose from a desire to obtain a common framework for an array of recent results dealing with Schatten class $S^{p}$ properties of commutators
$\phi\mapsto[f,T]\phi:=fT(\phi)-T(f\phi)$
of pointwise multipliers $f$ and singular integral operators $T$ in a variety of different settings [19, 20, 21, 26, 43] beyond the unweighted Euclidean space $\mathbb{R}^{d}$, where such results go back to [35, 50], as well as [15, Appendix] in a critical borderline case.
The proofs of such results naturally split into two parts of different spirit:
  1. (a)
    An interplay of operators and functions, where the Schatten norms of the commutator $[f,T]$ are connected to abstract oscillatory norms of the function $f$.
  2. (b)
    Analysis of functions only, where these abstract norms are further characterised in terms of more concrete and familiar function space properties.
The aim of this paper is to provide the relevant results of type (b) that, combined with results of type (a) in the companion paper [34], give a general characterisation of the Schatten class membership of commutators stated in [34, Theorem 1.1].
More specifically, in a general setting of a metric measure space $(X,\rho,\mu)$, our goal is to provide the essential function space theory underlying the following key phenomena in the Schatten class properties of commutators:
@@
    The cut-off phenomenon of [35] says that, for a threshold value $d$, we have $f,T]\in S^{p}$ for $p\leq d$ if and only if $f$ is constant (and hence $[f,T]=0$).
@@
    At the critical index $d$, the weak-type Schatten property $f,T]\in S^{d,\infty}$ is characterised (combining results of [15, 50]) by a first-order Sobolev norm of $f$.
Aside from this initial motivation and application, we believe that our results also have independent interest in the context of the following two parallel themes that have been extensively studied both in $\mathbb{R}^{d}$ and beyond:
  1. (1)
    Characterisation of constant functions by the finiteness of suitable integral expressions—a theme going back to [11] in $\mathbb{R}^{d}$ and subsequently extended to various settings in [7, 8, 30, 40, 48, 49].
  2. (2)
    Derivative-free characterisations of Sobolev spaces, several versions of which have been explored even in $\mathbb{R}^{d}$ [1, 9, 10, 12, 13, 15, 23, 44, 45, 46, 47], while extensions and variants over other spaces appear in [7, 16, 17, 29, 41, 42, 43, 52].
However, rather than just extending these lists, our present characterisations are specifically tailored for the needs of the commutator questions discussed above, and their usefulness is confirmed by their successful application in the companion work [34]. Our main theorem is the following generalisation of a recent Euclidean result of [23] and its Carnot group extension [43]; we will say more about the relations to earlier literature after the statement:
1.1 Theorem.
Let $(X,\rho,\mu)$ be a complete doubling metric measure space supporting a $(1,p)$-Poincaré inequality (Definition 2.4) for some $p\in(1,\infty)$. Let $f\in L^{1}_{\operatorname{loc}}(\mu)$. Then $f$ belongs to the homogeneous Hajłasz–Sobolev space $\dot{M}^{1,p}(\mu)$ (Definition 2.10) if and only if $m_{f}\in L^{p,\infty}(\nu_{p})$, where $m_{f}$ and $\nu_{p}$ are the function and the measure on $X\times(0,\infty)$ defined by
In this case, we have
$\|f\|_{\dot{M}^{1,p}(\mu)}$ $\sim\|m_{f}\|_{L^{p,\infty}(\nu_{p})}:=\sup_{\kappa>0}\kappa\cdot\nu_{p}(\{m_{f}>\kappa\})^{\frac{1}{p}}$ (1.3)
$\sim\liminf_{\kappa\to 0}\kappa\cdot\nu_{p}(\{m_{f}>\kappa\})^{\frac{1}{p}}.$ (1.4)
The norm comparison (1.3) is an exact analogue of the corresponding results in $\mathbb{R}^{d}$ [23, Theorem 1.1] and Carnot groups [43, Theorem A.3], but its proof will require substantial departures from [23, 43] in the lack of both group structure and higher-order smoothness. See Examples 2.12 and 2.14 for the coincidence of $\dot{M}^{1,p}(\mu)$ with the usual Sobolev spaces in the respective settings. On the other hand, our limit relation (1.4) is slightly weaker than the corresponding identity
$\|f\|_{\dot{W}^{1,p}(\mathbb{R}^{d})}=c_{d,p}\lim_{\kappa\to 0}\kappa\cdot\nu_{p}(\{m_{f}>\kappa\})^{\frac{1}{p}}$
proved in [23, Eq. (1.2)], where $c_{d,p}$ is an explicit constant, and the existence of the limit is part of the result. In the lack of the symmetries available in $\mathbb{R}^{d}$, we believe that (1.4) is the best substitute that one can expect in our generality.
Another previous result in the spirit of Theorem 1.1 is [7, Theorem 1.4]: in addition to our assumptions, it requires the space to be compact and Ahlfors $d$-regular, and only deals with integrability $p=d$ matching the homogeneous dimension of the space—in the spirit of [15], which imposed the same assumption on $\mathbb{R}^{d}$.
In contrast to the recent results of [16, 29] (which extend a different characterisation of the Sobolev space from [12, 13]), we would like to stress that Theorem 1.1 characterises the whole space $\dot{M}^{1,p}(\mu)$, rather than just giving an equivalent norm of test functions that are already known to be in the space. Indeed, some of the key challenges in the proof are about showing that the finiteness of the oscillatory norm $\|m_{f}\|_{L^{p,\infty}(\nu_{p})}$ ensures that $f\in\dot{M}^{1,p}(\mu)$, without a priori knowing whether $f$ belongs to this space.
Concerning the characterisation of constants, our second main result is the following, where we denote $V(x,y):=\mu(B(x,\rho(x,y)))$ for all $x,y\in X$.
1.5 Theorem.
Let $(X,\rho,\mu)$ be a doubling metric measure space. For $p,s\in(0,\infty)$, consider the following two variants of homogeneous Besov spaces:
$\begin{split}\dot{B}^{s}_{p,p}(\mu)&:=\Big\{f\in L^{p}_{\operatorname{loc}}(\mu):\iint_{X\times X}\Big(\frac{|f(x)-f(y)|}{\rho(x,y)^{s}}\Big)^{p}\frac{\,\mathrm{d}\mu(x)\,\mathrm{d}\mu(y)}{V(x,y)}<\infty\Big\},\\ \dot{B}^{p}(\mu)&:=\Big\{f\in L^{p}_{\operatorname{loc}}(\mu):\iint_{X\times X}|f(x)-f(y)|^{p}\frac{\,\mathrm{d}\mu(x)\,\mathrm{d}\mu(y)}{V(x,y)^{2}}<\infty\Big\}.\end{split}$
Let $d\in[1,\infty)$, and suppose that $(X,\rho,\mu)$ satisfies a $(1,d)$-Poincaré inequality.
  1. (1)
    Then $\dot{B}^{1}_{d,d}(\mu)=\{\operatorname{constants}\}$.
  2. (2)
    If, in addition, $(X,\rho,\mu)$ has lower dimension $d$ (Definition 2.1), then
    $\dot{B}^{p}(\mu)=\{\operatorname{constants}\}\qquad\forall\ p\in(0,d].$
We would like to stress that Theorem 1.5, in contrast to Theorem 1.1, does not assume the completeness of $X$. This reflects the relatively simpler considerations that enter into its proof.
It seems that $\dot{B}^{s}_{p,p}(\mu)$ is the most relevant version of a Besov (or fractional Sobolev) space over a metric measure space for many purposes [5, 6, 24, 25], but the variant $\dot{B}^{p}(\mu)$ is one that naturally arises in the context of commutators [20, 34]. If $X$ is Ahlfors $d$-regular, thus $V(x,y)\sim\rho(x,y)^{d}$, then $\dot{B}^{p}(\mu)=\dot{B}^{d/p}_{p,p}(\mu)$—consistent with the role of $\dot{B}^{d/p}_{p,p}(\mathbb{R}^{d})$ in classical commutator results of [35]. In this case, the condition $p\leq d$ means that $s=d/p\geq 1$, reflecting the fact that $\dot{B}^{s}_{p,p}$ above is not the right definition of the Besov space on $\mathbb{R}^{d}$ for smoothness $s\geq 1$.
Case (1) of Theorem 1.5 is recently due to [40, Corollary 1.8], but we provide a different approach of potentially independent interest, being relatively direct in using only the Besov spaces as defined above, while the proof of [40, Corollary 1.8]—quoting [2, Theorem 5.1], [32, Theorem 10.5.2], and [40, Theorem 1.6]—passes through the theory of Korevaar–Schoen spaces $KS^{1,p}$. In fact, we prove a slightly stronger local version of case (1) in Proposition 4.1.
On the other hand, case (2) is a new result to our knowledge, and it is this variant that is relevant for commutator applications, both in the companion paper [34] and in the recent [20]: In the specific “Bessel setting” $(X,\rho,\mu)=(\mathbb{R}_{+}^{n+1},|x-y|,x_{n+1}^{2\lambda}\,\mathrm{d}x)$ with $d=n+1$, [20, Proposition 6.4] obtain a version of case (2) in the weaker form
$C^{2}(\mathbb{R}_{+}^{n+1})\cap\dot{B}^{p}(x_{n+1}^{2\lambda}\,\mathrm{d}x)=\{\operatorname{constants}\}\qquad\forall\ p\in(0,n+1].$
The fact that our Theorem 1.5 leads to an improvement of this previous result, even when specialised to such a concrete situation, illustrates the power of our abstract methods. See Remark 4.7 below for further elaboration on this. Again, we also obtain a slightly stronger local variant of case (2) in Corollary 4.3.
Both Theorems 1.1 and 1.5 depend on the (by now usual, cf. [31, 32]) assumption of a Poincaré inequality, which we recall in Definition 2.4. Roughly speaking, this inequality says that oscillations of a function on a macroscopic scale are controlled by ones on the infinitesimal scale. As a key intermediate step between this assumption and Theorems 1.1 and 1.5, we establish a new “macroscopic Poincaré inequality”: oscillations on a larger macroscopic scale are controlled by ones on a smaller macroscopic scale, uniformly over the two scales. Formulated in Theorem 3.4, this is our third main result, which should have independent interest and other applications elsewhere.
1.A. Plan of the paper and an overview of the argument
In the following Section 2, we collect some preliminaries, including the definitions omitted in this Introduction. In Section 3, we present our key new tool of the “macroscopic” Poincaré inequality in Theorem 3.4. With this result, we obtain a rather quick proof of Theorem 1.5 on the characterisation of constant functions in Section 4.
The rest of the paper is dedicated to the proof of our Main Theorem 1.1 on the characterisation of the Sobolev space. The rough outline of our proof is adapted from the Euclidean argument of [23]. In particular, the lower bound of the Sobolev norm (i.e., $\gtrsim$” in (1.3)) is obtained in Section 5 by a straightforward adaptation of [23]. However, the other direction is much more challenging, and we need to develop substantial new methods. Besides the macroscopic Poincaré inequality, which also features in this proof, we establish some delicate comparison of maximal and mean oscillations at the infinitesimal scale in Section 6, to substitute tools like Taylor expansions, which are not available in our setting. However, we only obtain this comparison for Lipschitz functions. To lift this a priori assumption, we need to be able to approximate general functions by Lipschitz functions with respect to the “exotic” norm $f\mapsto\|m_{f}\|_{L^{p,\infty}(\nu_{p})}$ featuring in the desired characterisation. (This approximation follows easily from Theorem 1.1 a posteriori, once we know that the exotic norm is equivalent to $\|f\|_{\dot{M}^{1,p}(\mu)}$; but we need it a priori, in order to prove this equivalence!) This is accomplished in Section 7 with the help of the macroscopic Poincaré inequality, this time serving to substitute some convolution tricks, which are easy in $\mathbb{R}^{d}$ [23], and also on Carnot groups [43], but unavailable in our setting. After these preparations, we can complete the proof of the upper bound of the Sobolev norm ($\lesssim$” in (1.3)) in Section 8 and the limit relation (1.4) in Section 9.
In the final Section 10, we make some complementary remarks on different Poincaré inequalities. While not essential for our main results, these arise as natural by-products, and may have some independent interest.
2. Preliminaries
We will work over a metric measure space $(X,\rho,\mu)$. As usual, we write $A\lesssim B$, if $A\leq c\cdot B$ for some constant $c$ that only depends on the space $(X,\rho,\mu)$ and some related fixed parameters, but never on individual points $x\in X$, functions $f$, or similar variable quantities. We write $A\sim B$ if both $A\lesssim B$ and $B\lesssim A$.
@@ 2. Preliminaries
for all $x\in X$ and $r>0$. For $x,y\in X$, we also denote
$V(x,y):=V(x,\rho(x,y))\sim V(y,x),$
where the last equivalence is immediate from the doubling condition.
2.1 Definition.
A metric measure space $(X,\rho,\mu)$ is said to have lower dimension $d$ and upper dimension $D$ if
$\Big(\frac{R}{r}\Big)^{d}\lesssim\frac{V(x,R)}{V(x,r)}\lesssim\Big(\frac{R}{r}\Big)^{D}$
for all $x\in X$ and $0<r\leq R<2\operatorname{diam}(X)$. The space is said to be Ahlfors $d$-regular if it has lower and upper dimension $d=D$.
@@
Every doubling space has upper dimension $D\leq\log_{2}c_{\mu}$, but possibly much lower than this simple upper bound. While not a property of every doubling space, a lower dimension $d>0$ readily follows in the presence of a Poincaré inequality (Definition 2.4), which we assume in our main results: a space supporting a Poincaré inequality is connected by [4, Proposition 4.2], and a connected doubling space has lower dimension $d>0$ by [4, Corollary 3.8].
2.2 Remark.
The above definition of Ahlfors $d$-regular is equivalent to the usual one: a space if Ahlfors $d$-regular if and only if $V(x,r)\sim r^{d}$ for all $x\in X$ and $0<r<2\operatorname{diam}(X)$.
Proof.
“If” is clear. For “only if”, suppose that $V(x,R)/V(x,r)\sim(R/r)^{d}$ for all $x\in X$ and $0<r\leq R<2\operatorname{diam}(X)$. It is immediate that this implies doubling, and hence $V(x,R)\sim V(y,R)$ for $R\geq\rho(x,y)$. Given $x,y\in X$ and $0<r<2\operatorname{diam}(X)$, we can choose $R\in[r,2\operatorname{diam}(X))$ such that $R\sim r+\rho(x,y)$ to obtain
$V(x,r)\sim\Big(\frac{r}{R}\Big)^{d}V(x,R)\sim\Big(\frac{r}{R}\Big)^{d}V(y,R)\sim V(y,r).$
Fixing some $x_{0}\in X$, it follows that
$V(x,r)\sim V(x_{0},r)\sim\Big(\frac{r}{1}\Big)^{d}V(x_{0},1)\sim r^{d}.$